# Classical and quantum thermodynamics in a non–equilibrium regime: Application to Stirling engine

### v1

@@ Abstract
We develop a thermodynamic theory in the non–equilibrium regime by extending a theory based on the dimensionless (DL) minimum work principle previously developed for a thermodynamic system–bath model [S. Koyanagi and Y. Tanimura, J. Chem. Phys. in press (http://arxiv.org/abs/2405.16787)]. Our results are described by non–equilibrium thermodynamic potentials expressed in time-derivative form in terms of extensive and intensive variables. This is made possible through the incorporation of waste heat, which is equivalent to the loss work consumed by the bath, into the definitions of thermodynamic potentials in a non–equilibrium regime. These potentials can be evaluated from the DL non–equilibrium-to-equilibrium minimum work principle, which is derived from the principle of DL minimum work and is equivalent to the second law of thermodynamics. We thus obtain the non–equilibrium Massieu–Planck potentials as entropic potentials and the non–equilibrium Helmholtz–Gibbs potentials as free energies. Our results are verified numerically by simulating a Stirling engine consisting of two isothermal and two thermostatic processes using the quantum Fokker–Planck equations and the classical Kramers equation derived from the thermodynamic system–bath model. We then show that any thermodynamic process can be analyzed using a non–equilibrium work diagram analogous to the equilibrium one for given time-dependent intensive variables. The results can be used to develop efficient heat machines in non–equilibrium regimes.
@@ I Introduction
Ever since Carnot explored the efficiency of heat engines 200 years ago,Carnot (1824) there have been longstanding attempts to study thermodynamics in non–equilibrium regimes, driven by both academic curiosity and practical interest. In particular, recent advances in nanotechnology have led to increased interest in the study of the thermodynamics of microscopic quantum systems. Quan et al. (2006); Esposito, Harbola, and Mukamel (2009); Campisi, Hänggi, and Talkner (2011); Kosloff and Levy (2014); Esposito, Ochoa, and Galperin (2015); Schmidt et al. (2015); Uzdin, Levy, and Kosloff (2015); Whitney (2018); Quan et al. (2007); Maruyama, Nori, and Vedral (2009); Ono et al. (2020); Maruyama, Nori, and Vedral (2009); Talkner and Hänggi (2020a); Cangemi et al. (2020); Strasberg and Winter (2021); Cockrell and Ford (2022); Talkner and Hänggi (2020b); Dann and Kosloff (2023); Ferreri et al. (2023); Sakamoto and Tanimura (2021); Kato and Tanimura (2015, 2016)
@@
As a kinetic theory, such phenomena are typically explained using open quantum dynamics theories based on the system–bath (SB) model.Leggett et al. (1987); Caldeira and Leggett (1983); Grabert, Schramm, and Ingold (1988); Weiss (2012) Although thermodynamics is a system-independent theory, it is derived from the presence of a heat bath and is consistent with theories based on the SB model.Talkner and Hänggi (2020b) The main (subsystem) system of the SB model can describe a variety of systems, from a simple two-level system to complex molecular systems with many degrees of freedom, and it is also possible to take the classical limit in the case where the system can be described in phase space.Caldeira and Leggett (1983); Grabert, Schramm, and Ingold (1988); Weiss (2012); Tanimura and Kubo (1989); Tanimura (1990, 2014, 2006, 2020) Under the condition of a heat bath with infinite degrees of freedom, the total system irreversibly relaxes toward its equilibrium state in time, and so the second law of thermodynamics is naturally obeyed without the need for the assumption of ergodicity as a dynamical system.Tanimura (2006, 2020, 2014, 2015)
@@
However, to investigate whether the thermodynamic laws are satisfied even in quantum cases where the subsystem and bath are quantum mechanically entangled, it is necessary to treat the bath in a non-Markovian and nonperturbative manner so that the thermal equilibrium state of the total systems satisfies the energy conservation law, including the SB interaction. Therefore, equations of motion derived using the Markovian or rotational wave (or secular) approximation, such as the Lindblad equation and the quantum master equation, can only be applied to high-temperature regions where the subsystem exhibits semiclassical dynamics.Tanimura (2006, 2020, 2014, 2015) Over the past 30 years, several methodologies have been developed to accurately describe the effects of quantum entanglement for condensed systems while maintaining strict energy conservation to satisfy the first law of thermodynamics. Such methods include the hierarchical equations of motion (HEOM), Tanimura and Kubo (1989); Tanimura (1990); Ishizaki and Tanimura (2005); Tanimura (2006, 2020, 2014, 2015) the quasi-adiabatic path integral (QUAPI),Makri (1995); Makri and Makarov (1995a, b); Thorwart, Reimann, and Hänggi (2000); Jadhao and Makri (2008); Makri (2014); Segal, Millis, and Reichman (2010) and multiconfigurational time-dependent Hartree (MCTDH).Meyer, Manthe, and Cederbaum (1990); Manthe, Meyer, and Cederbaum (1992); Wang and Thoss (2007) Among these, the HEOM approach is ideal for thermodynamic investigations, not only because it can perform numerically “exact” dynamic simulations, but also because it can evaluate the change in energy of the heat bath and the SB interaction separately, even for processes far from equilibrium.Kato and Tanimura (2015, 2016); Sakamoto and Tanimura (2020, 2021); Song and Shi (2017); Gong et al. (2022); Latune, Pleasance, and Petruccione (2023); Boettcher et al. (2024); Koyanagi and Tanimura (2022a, b, 2024a)
@@
While the difficulty of numerical simulation has been solved, there is a fundamental difference between open quantum dynamics theory, which is based on a first-principles description of kinetic systems from a microscopic perspective, and thermodynamic theory, which is based on a phenomenological description of thermal systems from a macroscopic perspective. For example, in quantum mechanics, observables are defined as expectation values, whereas in thermodynamics, they are described by macroscopic intensive and extensive variables. Furthermore, quantum mechanics is a formalism for time evolution, whereas thermodynamics deals primarily with static and quasi-static states near thermal equilibrium. In this respect, most studies of quantum thermodynamics are merely kinetic simulations of open quantum dynamics, and their relation to thermodynamics has not been studied in depth. Thus, the fundamental difference between microscopic quantum mechanics and macroscopic thermodynamics raises many open questions, such as whether the Carnot limit can be violated in a quantum case or the existence of Maxwell’s demon.
@@
The virtue of thermodynamics lies in its ability to describe macroscopic thermal phenomena resulting from complex microscopic interactions in a system-independent manner as changes in thermodynamic potentials described as interrelated intensive and extensive variables through Legendre transformations. This virtue should be preserved when we develop a quantum thermodynamic theory rather than an open quantum dynamical theory, although in either case, the theory must be specific to the SB model.
@@
It has been found that the Gibbs energy can be obtained directly from kinetic simulations or experiments based on the minimum work principle (or Kelvin–Planck statement), expressed as $W(t)\geq\Delta G(t)$, where $W(t)$ is the work done from the outside to the subsystem by external fields and $\Delta G(t)$ is the change in free energy, by evaluating the work in a quasi-static process,Sakamoto and Tanimura (2020, 2021) which allows us to draw a work diagram corresponding to the $P$$V$ diagram.Koyanagi and Tanimura (2022a, b) However, with this approach, the contributions of temperature $T$ and entropy $S$ in the thermodynamic potential expressed as $TdS$ and $SdT$ cannot be evaluated, because the minimum work principle is defined as an isothermal process $dT=0$.
@@
To overcome this limitation, we developed a thermostatic SB model that is defined by a system coupled to multiple heat baths with different temperatures.Koyanagi and Tanimura (2024a) We then extended the minimum work principle to thermostatic processes in a dimensionless (DL) form (the DL minimum work principle) as ${\tilde{W}}(t)\geq\Delta\Xi(t)$, where ${\tilde{W}}(t)\equiv\beta(t)W(t)$ is the DL (entropic) work, with $\beta(t)\equiv 1/k_{\rm B}T(t)$ (where $k_{\rm B}$ is the Boltzmann constant) being the time-dependent inverse temperature, and $\Delta\Xi$ is the change in the DL Planck potential.Koyanagi and Tanimura (2024a) Not only intensive variables but also extensive variables, which are related by the time-dependent Legendre transformations, were introduced as quantum expectation values of the SB system. The validity of these results has been numerically verified using the low-temperature quantum Fokker–Planck equations in the quantum case and the Kramers equation in the classical case,Ikeda and Tanimura (2019) both developed for the thermodynamic SB model.Koyanagi and Tanimura (2024a) Although these results were restricted to quasi-static cases, the definitions of the thermodynamic potentials, intensive and extensive variables, and Legendre transformations were defined in such a way that they hold for any non–equilibrium process. Taking advantage of this, we extend our thermodynamic theory here to the non–equilibrium regime.
@@
The rest of this paper is organized as follows. In Sec. II, our previous results are summarized with regard to thermodynamics as applied to work in a system–independent manner. We then derive the principle of non–equilibrium DL minimum work to obtain the DL Massieu–Planck potentials and Helmholtz–Gibbs potentials in the non–equilibrium regime. Results are verified in Sec. III by numerical simulations using the SB model. Finally, Sec. IV presents concluding remarks.
II Reflections on work of heat
@@
In our previous paper,Koyanagi and Tanimura (2024a) we presented system–specific thermodynamic laws based on open quantum dynamics theory. Here, we develop the same laws by introducing several thermodynamic “statements” without going into the details of the system, as in traditional thermodynamic theory. By doing so, we clarify the distinctive features of thermodynamic theory that allow it to treat systems in non–equilibrium regimes.
II.1 Laws of thermodynamics as applied to work
We consider a thermodynamic system consisting of a subsystem A and a heat bath B at an inverse temperature $\beta(t)$. The presence of a heat bath temperature can be regarded as a consequence of the zeroth law of thermodynamics.
@@ II.1 Laws of thermodynamics as applied to work
An important statement of thermodynamics is that thermodynamic systems are described in terms of intensive and extensive variables. In particular, extensive variables that are proportional to the size are essential. This statement, controversial, is sometimes called the fourth law.Oono (2017) However, this is the premise of thermodynamics and we should call the minus first law of thermodynamics.
@@
The energy of the subsystem corresponds to internal energy and is expressed as $U_{\rm A}(t)$, which changes with time owing to changes in the bath temperature. Internal energy is an extensive variable, whereas inverse temperature is an intensive variable. The external perturbation we consider here is expressed as $-x(t)X_{\rm A}(t)$, where $x(t)$ and $X_{\rm A}(t)$ are intensive and extensive variables, respectively. In traditional thermodynamics, $x(t)$ is derived from the Euler relation. We then introduce the total energy, which is related to enthalpy by ${H}_{\rm A}(t)=U_{\rm A}(t)-x(t)X_{\rm A}(t)$, which can also be regarded as the Legendre transformation between internal energy and enthalpy.
@@
Because $U_{\rm A}(t)$ is the conjugate variable of $\beta(t)$, we introduce the DL (or entropic) internal energy denoted by $\tilde{U}_{\rm A}(t)$, where $\tilde{B}(t)\equiv\beta(t)B(t)$ for any variable $B(t)$. Because $\beta(t)$ diverges for $T\rightarrow 0$, states with $T=0$ do not exist, which corresponds to the third law of thermodynamics.
From the time derivative of the DL internal energy $\tilde{U}_{\rm A}(t)$, we have
@@
$\frac{d\tilde{Q}_{\rm A}^{ext}(t)}{dt}\equiv\beta(t)\frac{dU_{\rm A}(t)}{dt}-\tilde{x}(t)\frac{dX_{\rm A}(t)}{dt}$ (3)
are the DL (or entropic) extensive work defined as being done from the outside and DL (or entropic) extensive heat, respectively. For the enthalpy, we have
@@
is the intensive work. Equations (1) and (4) correspond to the first law of thermodynamics.
From the definitions, these intensive and extensive variables satisfy the time-dependent Legendre (TDL) transformation expressed as
@@
Equation (1) is also expressed in the form of the TDL transformation as
@@
The intensive heat $\tilde{Q}_{\rm A}^{int}(t)$ can also be defined from $\tilde{Q}_{\rm A}^{ext}(t)$ using the TDL [Eq. (46) in Appendix A]. While the above equations, including the first law of thermodynamics, hold for any non–equilibrium processes described by intensive variables $\beta(t)$ and $\tilde{x}(t)$ and extensive variables $U_{\rm A}(t)$ and $X_{\rm A}(t)$ which are the state variables, as they are kinematic observables at time $t$, work and heat are not state variables.
As the second law of thermodynamics, we adopt the DL minimum work principle for the subsystem from one equilibrium state to another, which extends the Kelvin-Planck statement for isothermal processes to thermostatic processes. For work defined as being done from the outside to the subsystem, we have
@@
where $\Phi_{\rm A}^{\rm qst}$ is the DL Massieu potential Massieu (1869) and equality holds under quasi-static changes of natural variables $\beta(t)$ and $X_{\rm A}(t)$ as $\Phi_{\rm A}^{\rm qst}[\beta^{\rm qst},X_{\rm A}^{\rm qst}]=\tilde{W}_{\rm A}^{ext}[\beta^{\rm qst},X_{\rm A}^{\rm qst}]$. The above inequality states that for any process occurring between two states, the work done from the outside is minimized if the process is quasistatic (or reversible), because no energy is dissipated in the heat bath. Note that if we define work as being done from the subsystem to outside, the inequality in Eq.(8) is reversed and the relationship is called the maximum work principle (to the outside). Using the TDL transformation (6), we also have the Planck potential Planck (1922) in terms of natural variables expressed as $\Xi_{\rm A}^{\rm qst}[\beta^{\rm qst},\tilde{x}^{\rm qst}]$ (see Appendix A).

### v2

# Classical and quantum thermodynamics in a non-equilibrium regime: Application to thermostatic Stirling engine

@@ Abstract
We have developed a thermodynamic theory in the non-equilibrium regime, which we describe as a thermodynamic system-bath model [S. Koyanagi and Y. Tanimura, J. Chem. Phys. 160, 234112 (2024)]. Based on the dimensionless (DL) minimum work principle, non-equilibrium thermodynamic potentials are expressed in terms of non-equilibrium extensive and intensive variables in time derivative form. This is made possible by incorporating the entropy production rate into the definitions of non-equilibrium thermodynamic potentials. These potentials can be evaluated from the DL non-equilibrium-to-equilibrium minimum work principle, which is derived from the principle of DL minimum work and is equivalent to the second law of thermodynamics. We thus obtain the non-equilibrium Massieu–Planck potentials as entropic potentials and the non-equilibrium Helmholtz–Gibbs potentials as free energies. Unlike fluctuation theorem and stochastic thermodynamics theory, this theory does not require the assumption of a factorized initial condition and is valid in the full quantum regime where the system and bath are quantum mechanically entangled. Our results are numerically verified by simulating a thermostatic Stirling engine consisting of two isothermal processes and two thermostatic processes using the quantum hierarchical Fokker–Planck equations and the classical Kramers equation derived from the thermodynamic system-bath model. We then show that, from weak to strong system-bath interactions, the thermodynamic process can be analyzed using a non-equilibrium work diagram analogous to the equilibrium one for given time-dependent intensive variables. The results can be used to develop efficient heat machines in non-equilibrium regimes.
@@ I Introduction
Ever since Carnot explored the efficiency of heat engines 200 years ago,Carnot (1824) there have been longstanding attempts to study thermodynamics in non-equilibrium regimes driven by academic curiosity and practical interest. In particular, recent advances in nanotechnology have led to increased interest in the study of the thermodynamics of microscopic systems. Quan et al. (2006); Kosloff and Levy (2014); Esposito, Ochoa, and Galperin (2015); Schmidt et al. (2015); Uzdin, Levy, and Kosloff (2015); Whitney (2018); Quan et al. (2007); Maruyama, Nori, and Vedral (2009); Ono et al. (2020); Maruyama, Nori, and Vedral (2009); Cangemi et al. (2020); Strasberg and Winter (2021); Esposito, Harbola, and Mukamel (2009); Campisi, Hänggi, and Talkner (2011); Sagawa (2023); Schmiedl and Seifert (2007); Seifert (2012); Esposito, Lindenberg, and den Broeck (2010); Esposito and den Broeck (2011); Soret, Cavina, and Esposito (2022); Strasberg (2022); Rao and Esposito (2018); Cockrell and Ford (2022); Talkner and Hänggi (2020); Dann and Kosloff (2023); Usui et al. (2024); Soret, Cavina, and Esposito (2022); Ferreri et al. (2023); Sakamoto and Tanimura (2020, 2021); Kato and Tanimura (2015, 2016)
@@
As a kinetic theory, such phenomena are typically explained using open quantum dynamics theories based on the system–bath (SB) model.Leggett et al. (1987); Caldeira and Leggett (1983); Grabert, Schramm, and Ingold (1988); Weiss (2012) Although thermodynamics is a system-independent theory, it is derived from the presence of a heat bath and is consistent with theories based on the SB model.Talkner and Hänggi (2020) The main (subsystem) system of the SB model can be described from a simple two-level system to complex molecular systems with many degrees of freedom, and it is also possible to take the classical limit in the case of a system described in phase space.Caldeira and Leggett (1983); Grabert, Schramm, and Ingold (1988); Weiss (2012); Tanimura and Kubo (1989); Tanimura (1990, 2014, 2006, 2020) Under the condition of a heat bath with infinite degrees of freedom, the total system irreversibly relaxes toward its equilibrium state in time, so the second law of thermodynamics is naturally obeyed without the need for the assumption of ergodicity as a dynamical system.Tanimura (2006, 2020, 2014, 2015)
However, to investigate whether the thermodynamic laws are satisfied even in quantum cases where the subsystem and bath are quantum mechanically entangled, it is necessary to treat the bath in a non-Markovian, non-perturbative, and non-factorized manner so that the thermal equilibrium state of the total system satisfies the energy conservation law, including the SB interaction. Therefore, equations of motion derived using the Markovian or rotational wave (or secular) approximation, such as the Lindblad equation and quantum master equation, can only be applied to high-temperature regions where the subsystem exhibits semiclassical dynamics.Tanimura (2006, 2020, 2014, 2015) Therefore, the validity of theories based on the Markov assumptionVan Vu and Saito (2022) or those based on factorized initial conditions, such as the fluctuation theoremEsposito, Harbola, and Mukamel (2009); Campisi, Hänggi, and Talkner (2011); Sagawa (2023) and stochastic thermodynamicsSchmiedl and Seifert (2007); Seifert (2012); Esposito, Lindenberg, and den Broeck (2010); Esposito and den Broeck (2011); Strasberg (2022); Rao and Esposito (2018); Cockrell and Ford (2022); Soret, Cavina, and Esposito (2022) in the quantum case, should be carefully examined.
@@
Over the past 30 years, several methodologies have been developed to accurately describe the effects of quantum entanglement in condensed systems while maintaining strict energy conservation to satisfy the first law of thermodynamics. Such methods, in order of development, include the hierarchical equations of motion (HEOM), Tanimura and Kubo (1989); Tanimura (1990); Ishizaki and Tanimura (2005); Tanimura (2006, 2020, 2014, 2015) the quasi-adiabatic path integral (QUAPI),Makri (1995); Makri and Makarov (1995a, b); Thorwart, Reimann, and Hänggi (2000); Jadhao and Makri (2008); Makri (2014); Segal, Millis, and Reichman (2010) and multiconfigurational time-dependent Hartree (MCTDH),Meyer, Manthe, and Cederbaum (1990); Manthe, Meyer, and Cederbaum (1992); Wang and Thoss (2007) in historical order. Among these, the HEOM approach is ideal for thermodynamic investigations, not only because it can perform numerically “exact” dynamic simulations but also because it can evaluate the change in energy of the heat bath and the SB interaction separately, even for processes far from equilibrium.Kato and Tanimura (2015, 2016); Sakamoto and Tanimura (2020, 2021); Song and Shi (2017); Gong et al. (2022); Latune, Pleasance, and Petruccione (2023); Boettcher et al. (2024); Koyanagi and Tanimura (2022a, b, 2024a, 2024b, 2024c) Note that although not numerically “exact”, several non-perturbative approaches have also been developed specifically for thermodynamic systems.Strasberg et al. (2018); Strasberg and Esposito (2020); Zhou, Li, and Galperin (2024)
@@
Although the difficulty of numerical simulation has been solved, a fundamental difference exists between open quantum dynamics theory, which is based on a first-principles description of kinetic systems from a microscopic perspective, and thermodynamic theory, which is based on a phenomenological description of thermal systems from a macroscopic perspective. For example, in quantum mechanics, observables are defined as expectation values, whereas in thermodynamics, they are described by macroscopic intensive and extensive variables. Furthermore, quantum mechanics is a formalism for time evolution, whereas thermodynamics deals primarily with static and quasi-static states near thermal equilibrium. In this respect, most quantum thermodynamics studies are merely kinetic simulations of open quantum dynamics, and their relation to thermodynamics has not been studied in depth. Thus, the fundamental difference between microscopic quantum mechanics and macroscopic thermodynamics raises many open questions, such as whether the Carnot limit can be violated in a quantum case or the existence of Maxwell’s demon.
@@
The virtue of thermodynamics lies in its ability to describe macroscopic thermal phenomena resulting from complex microscopic interactions in a system-independent manner, as changes in thermodynamic potentials described as interrelated intensive and extensive variables through Legendre transformations. This virtue should be preserved when developing quantum thermodynamic theory rather than open quantum dynamical theory, although in either case, the theory must be specific to the SB model.
@@
The assumption of a factorized initial state is essential for the application of stochastic thermodynamics and the fluctuation theorem, so that these theories cannot describe the transitions between states in which the system and the bath are entangled. Also, in practice many quantum thermodynamics arguments treat a heat bath perturbatively, assuming Markovian time evolution,Esposito, Harbola, and Mukamel (2009); Campisi, Hänggi, and Talkner (2011); Sagawa (2023); Schmiedl and Seifert (2007); Seifert (2012) it has been found that the Gibbs energy can be obtained directly from kinetic simulations even in the case of non-Markovian and non-perturbative SB interaction at low temperature, where quantum entanglement between the system and the bath plays an essential role.Sakamoto and Tanimura (2020, 2021); Koyanagi and Tanimura (2022a, b) This approach based on the minimum work principle (or Kelvin–Planck statement), expressed as $W(t)\geq\Delta G(t)$, where $W(t)$ is the work done by the outside on the subsystem by external fields and $\Delta G(t)$ is the change in free energy, by evaluating the work in a quasi-static process,Sakamoto and Tanimura (2020, 2021) which allows us to draw a work diagram corresponding to the $P$$V$ diagram.Koyanagi and Tanimura (2022a, b) However, even with this approach, the contributions of temperature $T$ and entropy $S$ to the thermodynamic potential expressed as $TdS$ and $SdT$ cannot be evaluated because the minimum work principle is defined as an isothermal process $dT=0$.
@@
To overcome this limitation, we developed a thermostatic SB model that is defined by a system coupled to multiple heat baths at different temperatures.Koyanagi and Tanimura (2024b, c) We then extended the minimum work principle to thermostatic processes in a dimensionless (DL) form (the DL minimum work principle) as ${\tilde{W}}(t)\geq\Delta\Xi(t)$, where ${\tilde{W}}(t)\equiv\beta(t)W(t)$ is the DL (entropic) work, $\beta(t)\equiv 1/k_{\rm B}T(t)$ (where $k_{\rm B}$ is the Boltzmann constant) is the time-dependent inverse temperature, and $\Delta\Xi$ is the change in the DL Planck potential.Koyanagi and Tanimura (2024b) Not only intensive variables but also extensive variables, which are related by time-dependent Legendre transformations, were introduced as quantum expectation values of the SB system.
The validity of these results was verified using the numerically “exact” HEOM formalism. For thermodynamic studies, the HEOM approach has been used for spin-boson-based systems, taking advantage of the ability to evaluate the energy changes of the system, interaction, and bath, respectively, even under non-perturbative and non-Markovian conditions.Kato and Tanimura (2015, 2016); Sakamoto and Tanimura (2020, 2021); Koyanagi and Tanimura (2022a, b, 2024a) Here, we employed the HEOM formalism for an anharmonic quantum Brownian model to construct a thermodynamic theory that is valid for both classical and quantum cases. Restricting to the case of an Ohmic spectral distribution function (SDF), we derived the thermostatic quantum Fokker–Planck equations (T-QFPE)Koyanagi and Tanimura (2024b, c) on the basis of the low-temperature quantum Fokker–Planck equations (LT-QFPE) in the quantum case and the Kramers equation in the classical case.Ikeda and Tanimura (2019) In the classical and high-temperature limits, the T-QFPE are equivalent to the Langevin equation, where a Markovian description can be applicable, but at low temperatures, owing to quantum entanglement with the bath (bathentanglement),Tanimura (2020) the subsystem follows non-factorial and non-Markovian dynamics, and its equilibrium state deviates from the Boltzmann distribution. This indicates that thermodynamics in the fully quantum regime cannot be described by a theory based on the Markovian assumption.Koyanagi and Tanimura (2024c, a)
Although these results were restricted to quasi-static cases, the thermodynamic potentials, intensive and extensive variables, and Legendre transformations were defined in such a way that they hold for any non-equilibrium process. Taking advantage of this, we extend our thermodynamic theory here to the non-equilibrium regime.
@@
The remainder of this paper is organized as follows. In Sec. II, our previous results about thermodynamics as applied to work in a system-independent manner are summarized. We then derive the principle of non-equilibrium DL minimum work to obtain the DL Massieu–Planck potential and Helmholtz–Gibbs potentials in the non-equilibrium regime. Results are verified in Sec. III by numerical simulations using the thermostatic SB model. Finally, Sec. IV presents concluding remarks.
II Reflections on motive power of heat
@@
In our previous paper,Koyanagi and Tanimura (2024b) we presented system-specific thermodynamic laws described as a system–bath model on the basis of open quantum dynamics theory. Here, we develop the same laws by introducing several thermodynamic “statements” without going into the details of the system, as in traditional thermodynamic theory. By doing so, we clarify the distinctive features of thermodynamic theory that allow it to treat systems in non-equilibrium regimes.
II.1 Laws of thermodynamics applied to work
We consider a thermodynamic system consisting of subsystem A and heat bath B at the inverse temperature $\beta(t)$. The presence of a heat bath temperature can be regarded as a consequence of the zeroth law of thermodynamics, which states the existence of a unique equilibrium state.Oono (2017) Although not well known, the minus first law of thermodynamics was introduced some time ago to state the existence of time-irreversible processes toward a unique equilibrium state.Brown and Uffink (2001); Uffink (2001) The dynamics of the SB model comply with this law.
@@ II.1 Laws of thermodynamics applied to work
An important statement of thermodynamics is that thermodynamic systems are described in terms of intensive and extensive variables. In particular, extensive variables that are proportional to the size are essential. This statement is called the fourth law.Oono (2017) However, this is the premise of thermodynamics, and we would like to call it the minus second law of thermodynamics because this is the central dogma in the classical and quantum thermodynamics that we are constructing.
@@
The energy of the subsystem corresponds to the internal energy and is expressed as $U_{\rm A}(t)$, which changes with time owing to changes in the bath temperature. Internal energy is an extensive variable, whereas inverse temperature is an intensive variable. The external perturbation considered here is expressed as $-x(t)X_{\rm A}(t)$, where $x(t)$ and $X_{\rm A}(t)$ are intensive and extensive variables, respectively. In traditional thermodynamics, $x(t)$ is derived from the Euler relation. We then introduce the total energy, which is related to the enthalpy by ${H}_{\rm A}(t)=U_{\rm A}(t)-x(t)X_{\rm A}(t)$, which can also be regarded as the Legendre transformation between internal energy and enthalpy.
@@
Because $U_{\rm A}(t)$ is the conjugate variable of $\beta(t)$, we introduce the DL (or entropic) internal energy denoted by $\tilde{U}_{\rm A}(t)$, where $\tilde{B}(t)\equiv\beta(t)B(t)$ for any variable $B(t)$. Because $\beta(t)$ diverges for $T\rightarrow 0$, states with $T=0$ do not exist. Thus the introduction of DL variables corresponds to the third law of thermodynamics.
From the time derivative of the DL internal energy $\tilde{U}_{\rm A}(t)$, we obtain
@@
$\frac{d\tilde{Q}_{\rm A}^{ext}(t)}{dt}\equiv\beta(t)\frac{dU_{\rm A}(t)}{dt}-\tilde{x}(t)\frac{dX_{\rm A}(t)}{dt},$ (3)
and $\tilde{W}^{ext}_{\rm A}(t)$ and $\tilde{Q}_{\rm A}^{ext}(t)$ represent the DL (or entropic) extensive work defined as being done by the outside and DL (or entropic) extensive heat, respectively. The time derivative of $\tilde{W}^{ext}_{\rm A}(t)$ corresponds to the DL extensive power. For the enthalpy, we have
@@
and $\tilde{W}^{int}_{\rm A}(t)$ represents the DL intensive work. The time derivative of $\tilde{W}^{int}_{\rm A}(t)$ then corresponds to the DL intensive power. Equations (1) and (4) correspond to the first law of thermodynamics.
From the definitions, these intensive and extensive variables satisfy the time-dependent Legendre (TDL) transformation expressed as follows:
@@
Equation (1) can also be expressed in the form of a TDL transformation as
@@
The DL intensive heat $\tilde{Q}_{\rm A}^{int}(t)$ can also be defined from $\tilde{Q}_{\rm A}^{ext}(t)$ using the TDL [Eq. (46) in Appendix A]. The above equations, including the first law of thermodynamics, are described in terms of intensive variables $\beta(t)$ and $\tilde{x}(t)$ and extensive variables $U_{\rm A}(t)$ and $X_{\rm A}(t)$ and hold for any non-equilibrium processes. While these intensive and extensive variables are the state variables because they are kinetic observables at time $t$, work and heat are not state variables.
As the second law of thermodynamics, we adopt the DL minimum work principle for the subsystem moving from one equilibrium state to another, thereby extending the Kelvin–Planck statement for isothermal processes to thermostatic processes. For work defined as being performed from the outside on the subsystem, we have
@@
where $\Phi_{\rm A}^{\rm qst}$ is the DL Massieu potential Massieu (1869); Callen and equality holds under quasi-static changes in the natural variables $\beta(t)$ and $X_{\rm A}(t)$ as $\Phi_{\rm A}^{\rm qst}[\beta^{\rm qst},X_{\rm A}^{\rm qst}]=\tilde{W}_{\rm A}^{ext}[\beta^{\rm qst},X_{\rm A}^{\rm qst}]$. The above inequality states that for any process occurring between two states, the work done by the outside is minimized if the process is quasi-static (or reversible) because no energy is dissipated in the heat bath. Note that if we define work as being done by the subsystem on the outside, the inequality in Eq.(8) is reversed, and the relationship is called the maximum work principle (on the outside). Although not mathematically rigorous, the derivation of Eq. (8) is described in Appendix A of Ref. Koyanagi and Tanimura, 2024b. Using the TDL transformation (6), we also obtain the Planck potential Planck (1922); Callen in terms of natural variables expressed as $\Xi_{\rm A}^{\rm qst}[\beta^{\rm qst},\tilde{x}^{\rm qst}]$ (see Appendix A). Similar expressions for Planck functions were derived on the basis of stochastic thermodynamics.Esposito and den Broeck (2011); Rao and Esposito (2018); Strasberg (2022) However, our expression was obtained from the well-defined kinetic system and is valid even under non-factorized SB interactions, where bathentanglement plays an essential role. Thus, the value of $\Delta\Phi_{\rm A}^{\rm qst}$ appearing in Eq. (8) is generally different from the value evaluated from stochastic thermodynamics.